### v1

# On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems

Abstract
Large Language Models (LLMs) have been observed to perform well on a wide range of downstream tasks when fine-tuned on domain-specific data. However, such data may not be readily available in many applications, motivating zero-shot or few-shot approaches using domain-adjacent models. While several fine-tuned models for various tasks are available, finding an appropriate domain-adjacent model for a given task is often not straight forward. In this paper, we study DAFT-E, a framework that utilizes an Ensemble of Domain-Adjacent Fine-Tuned Foundation Models for few-shot problems. We show that for zero-shot problems, this ensembling method provides an accuracy performance close to that of the single best model. With few-shot problems, this performance improves further, at which point DAFT-E can outperform any single domain-adjacent model while requiring much less data for domain-specific fine-tuning.
On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems
Md Ibrahim Ibne Alam and Koushik Kar Rensselaer Polytechnic Institute NY, USA, 12180                        Parikshit Ram, Soham Dan and Horst Samulowitz IBM, Yorktown Heights NY, USA, 10598
1 Introduction
Pre-trained Large Language Models (LLMs), are used for different downstream tasks. Usually, for LLMs that are not directly suitable for downstream tasks, a task-specific header layer is needed at the output. We define the base model as the LLM with a task-specific header layer. As an example, BERT models Kenton and Toutanova (2019) were trained for sentence completion tasks. In order to perform sentiment classification, we need to add a header layer with output dimension matching the number of classes. Usually, base models do not need much training to perform well, because the pre-trained LLMs have already been trained on a large corpus of data. To perform a task with base models, we can fine-tune a base model with task-specific training data, and then use the fine-tuned model. However, it is often the case that appropriate training data for fine-tuning is not readily or abundantly available. Moreover, even if one does have the data to fine-tune, one may not have the computational resources, or the time required, to fine-tune these base models. It is worth noting that fine-tuning can be computation, memory and time intensive, even if very few iterations are needed to attain good performance. Thus, we explore alternatives that bypass those challenges and leverage the large number of fine-tuned LLMs that are already available in curated repositories such as Hugging-face HF (2024c); Kim (2023). It is important to acknowledge the scale of fine-tuned models being created. For instance, just three days after LLaMA-3 was released, more than 1,000 fine-tuned LLMs were publicly available on Hugging-face HF (2024a).
Here we show the potential of these existing and publicly available fine-tuned models (e.g., including those fine-tuned via LORA Sheng et al. (2023) and PEFT Li and Liang (2021)) by leveraging them to perform different tasks. To use a fine-tuned model on a task, e.g., sentiment classification, we need that model to be domain adjacent (formally defined and discussed in Section 3). Informally, a fine-tuned LLM is domain-adjacent if (i) it has been fine-tuned on a similar general task, such as sentiment analysis, and (ii) its output space can be appropriately mapped to the current (few-shot) task. We denote these Domain Adjacent Fine-Tuned models as DAFT. In the generation process of DAFT, the closer the fine-tuning dataset is to the target domain data, the better will be the performance of DAFT. One main property of DAFT is that it can be used without any further training (no extra computation) in a zero-shot manner (no data adaptation). To perform inference with DAFT we do not require large computational resources compared to when performing fine-tuning. It is important to note that when using DAFT, a method to choose among available model(s) to achieve improved performance for a given task is required (see Section 4).
The availability of DAFT models combined with their ease of use, make them very appealing to use for different tasks. To analyze the efficiency of using these DAFT models in comparison to other existing options, we first define several comparable alternatives. Let us denote $D_{T}$ as the testing dataset on which we need to perform a given task. Then we have the following options (Figure 1):
  1. 1.
    FT – Fine-tune base models with training data obtained from $D_{T}$.
  2. 2.
    DAFT ${}^{\text{Z}}$ – Pick one available DAFT model and use it for zero-shot inference (no training cost).
  3. 3.
    DA(FT)2 – Pick one available DAFT model and fine-tune with training data from $D_{T}$.
  4. 4.
    DAFT-E ${}^{\text{Z}}$ – Use a subset or all DAFT models and ensemble them for zero-shot inference.
  5. 5.
    DAFT-E – Use a subset or all DAFT models and perform few-shot learning of ensemble weights using training data from $D_{T}$ for inference.
Figure 1: Different options for few-shot tasks given LLMs and community created DAFT models: (1) Directly fine-tune the LLM as FT; (2) select one DAFT model and use it zero-shot as DAFT ${}^{\text{Z}}$, or (3) with task-specific fine-tuning as DA(FT)2. (4) DAFT models can be ensembled and used zero-shot as DAFT-E ${}^{\text{Z}}$, or (5) with extremely lightweight few-shot adaptation using weighted ensembling as DAFT-E.
In the next section we discuss related work, followed by Section 3 defining DAFT models and evaluating their performance on various NLP tasks. Section 4 discusses zero-shot and few-shot performance, while Section 5 provides some theoretical performance bounds for the ensembling methods.
2 Related Work
Ensembling is a technique that combines the predictions of multiple classifiers to generate a single decision and is extensively investigated Breiman (1996); Clemen (1989); Maclin and Opitz (1997); Dietterich (2000). The pre-requisites for effective ensembling are: (i) models with decent performance, and (ii) models that make independent errors, i.e., training diversity in models Ovadia et al. (2019); Lakshminarayanan et al. (2017). The ensemble of models can be the weighted or unweighted average of the model outputs. For weighted averaging, some training is needed to learn the weights Caruana et al. (2004). Isotropic merging Polyak and Juditsky (1992) is another technique similar to ensembling that has been studied alongside ensemble methods over the last few decades and has shown promising performance Wortsman et al. (2022). The main novelty of our work is the use of ensembling in the context of DAFT models fine-tuned from pretrained LLMs, to maximize the efficacy of using these models.
Another approach related to ensembling is blending of models. For instance, LLM-Blender Jiang et al. (2023) performs ensembling of $n$ LLMs by pairwise ranking and generative fusion. The pairwise ranking approach requires creating a custom dataset, training a BERT model, and it incurs substantial computational overhead when performing inference. The proposed generative fusion combines the ranked list from the pairwise ranking approach with a fine-tuned LLM for this task to generate a response. In contrast, our approaches focus on a completely task agnostic (e.g., not just generative tasks) and computationally low-cost solution to leverage multiple fine-tuned LLMs. Another type of blending is introduced in Lu et al. (2024), and it is orthogonal to our approach since it selects base models at random and combines them by adding the response of an already evaluated model to the input of the subsequent model. While this knowledge sharing makes it a sequential approach by design, it can be readily combined with ours.
Mixture of Experts (MoE) Jacobs et al. (1991) has been utilized by different ML models for both regression and classification tasks Yuksel et al. (2012). Before recent developments in LLMs, MoE was mainly focused on dense expert models. However, that idea has been transformed into sparse expert models with the introduction of MoE in LLMs. In current practice, MoE is a sparse expert model, where a set of parameters of a neural network are partitioned into experts having their own weights Fedus et al. (2022). One recent work on MoE, Mixtral Jiang et al. (2024) uses a sparse LLM MoE where a routing network (e.g., Rosenbaum et al. (2017)) is pre-trained to map input tokens to a subset of experts. Compared to all these MoE models, our approach does not require intensive pre-training, and the most computationally expensive variant that uses weight-learning (DAFT-E) is only trained in a few-shot fashion based on input examples.
Transfer learning Pan and Yang (2009); Zhuang et al. (2020) is closely related to the goals of our work here, where we wish to adapt a model trained on a data-rich source distribution to a data-scarce target problem. The success on the target problem often relies heavily on the level of positive transfer one can achieve, and various ways of training the model on the source data and the target data have been developed to maximize the positive transfer. In the context of LLMs, fine-tuning a (base) LLM with task-specific data is an instance of transfer learning. In our paper, we consider the DAFT models as the source models. However, not every DAFT model may transfer positively, since the DAFT models are obtained from public non-curated pools.
3 DAFT Models
With the rapid expansion of open platforms like Colab and Kaggle to perform model training, and having access to different public datasets, researchers can access LLMs and datasets that can be used to fine-tune these LLMs. This has lead to a large repository of publicly available fine-tuned base models (as defined in Section 1) HF (2024c); Kim (2023). In Section 1 we argued that the performance of any such fine-tuned model on some target dataset depends on the source dataset on which it was fine-tuned on. We define a dataset to be domain adjacent dataset if it can be used for the same (general) task as the target dataset. We call a model a domain adjacent fine-tuned model (DAFT) if: (i) it was fine-tuned with a domain adjacent dataset, and (ii) the output space of the model can be appropriately mapped to the target task.
3.1 Formation of DAFT models
3.1.1 Datasets
We consider two broad NLP tasks with datasets from different sources HF (2024b); University (2024); Kaggle (2024), for (1) Sentiment Analysis (positive or negative sentiment): Amazon polarity, Cornell Movie, IMDB, SST2, Tweet sentiment, and Yelp Polarity; and (2) Textual Similarity (if two sentences are similar): MRPC, QQP, STS-B (details in Appendix A.1).
3.1.2 Models
For both tasks, we chose three LLMs (with fine-tuning) for different performance analyses: (1) Roberta-base, (2) BERT-based-uncased, and (3) xlnet-base-cased. These models with a header layer form base models for our experiments. [^1] These three models were chosen from a few other models due to their usually better performance (on the chosen datasets) compared to all the other models of the same size. We also used Bart-large-mnli, Roberta-large-mnli, and Opt-1.3b models for sentiment analysis (zero-shot classification). These models are much larger in size compared to the above-mentioned models, and therefore needs more computational power for inference.
3.1.3 Base Models
A base model is created by adding a header layer to an LLM. Since the base models are not trained to perform any specific task and the header layer has not been fine-tuned, the performance of all the base models is similar to random guessing. For a specific dataset we can fine-tune the base models, and depending on the extent of fine-tuning, we obtain a partial (few-shot) or fully fine-tuned model. For some base model $B$ and dataset $D$, let us denote $\Phi(B,D,n)$ as the fine-tuned version of $B$ on $n$ amount of data from dataset $D$. If the parameter $n$ is missing as the argument, the base model is assumed to be fully fine-tuned (FFT) on $D$. In our experiments, with few shot fine-tuning, we vary $n$ in the range of $2-256$ samples. For the (FFT) models, we fine-tune until loss stabilization [^2] We fine-tune until the loss fluctuation of less than ±$\pm 1\%$..
3.2 Performance of DAFT Models
Figure 2: Efficacy of DAFT models: The vertical axis corresponds to the data used to fine-tune a base LLM (here we show Roberta-base; the heatmaps for the other models show a similar trend), and the horizontal axis corresponds to the unseen test data. All datasets have matching output spaces and pertain to a sentiment classification problem. The diagonal is the performance of FT– the LLM tested with data from the distribution it was fine-tuned with , denoting the upper watermark for that test data (only exception being SST2 – discussed in text). The off-diagonal entries correspond to the performance of the DAFT– LLM trained on one dataset, and tested on another, highlighting the potential computationally cheap (free!) zero-shot benefit that DAFT models can provide. Note that, for any given test data, the performance of a DAFT LLM can vary significantly.
Let us denote the train and test splits of the target dataset as $D_{T}^{\prime}$ and $D_{T}^{\prime\prime}$ respectively. Since the standard train and test splits are identically distributed (iid), we expect a model trained on the target data ($D_{T}^{\prime}$) to perform the best on the test data $D_{T}^{\prime\prime}$. This performance value is the ceiling benchmark that we aim to match or surpass. For that purpose, we performed full fine-tuning (FFT) on all three base models using training data of all datasets of sentiment analysis and textual similarity, to come up with 18 FFT models for sentiment analysis and 9 FFT models for textual similarity. Hence, for any target dataset, we can have 15 DAFT models for sentiment analysis by excluding the three models that were fully fine-tuned on $D_{T}$, and 6 DAFT models for textual similarity task.
We plot the individual heat maps with the FFT models on sentiment analysis and textual similarity tasks in Figure 2. It is evident from the figure that the performance of the Cornell and Tweet datasets do not match the other four datasets’ performance on the sentiment analysis task. Our empirical study suggests that the test data of Cornell is distributionally different from the other datasets, and for all FFT models the performance was poor. The Tweet dataset seems to have a different distribution in both its training and testing data (compared to the other 5 datasets). Looking at the Tweet row and column, we observe that apart from the diagonal element, all the other accuracies are small, and a high $94\%$ accuracy is attained when the model is trained and tested on Tweet data. Upon inspecting the row of SST2, another interesting observation can be made, the performance of DAFT-Cornell beats that of DAFT-SST2. This is caused by the test dataset of SST2 being more aligned with the training data of Cornell, with further details in Appendix A.3.1. Lastly, for textual similarity datasets, the performance of DAFT models was not as good as the sentiment analysis, but still show promise.
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]These results show that DAFT models can potentially provide very competitive zero-shot performance on new tasks.
3.2.1 Performance of DAFT ${}^{\text{Z}}$
Figure 3: Relative Improvement of DAFTs compared to the Single Best DAFT for sentiment analysis: For each test dataset, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and show the RI of DAFT ${}^{\text{Z}}$ over the single-best DAFT LLM (out of the 15) (the colored circles $\bullet$). Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot prompting on larger LLMs (Bart-large-mnli, Roberta-large-mnli, and Opt-1.3b).
Figure 3 shows the relative improvement $(RI)$ of all the DAFT models for sentiment analysis task when compared to the single best performing DAFT model. [^3] Thus, all the $RI$ values (corresponding to the other DAFTs) will be non-positive (zero for the best single DAFT). The relative improvement $(RI)$ of all the DAFT models for textual similarity task is provided in Appendix (Figure 14). From Figure 3 it is evident that there are some DAFT models that are not as good as the single best and for some specific choices (as for Tweet), the performance can be poor. On the other hand, among the three larger models (models from $\mathcal{L}$), one of them (opt-1.3b) showed poor performance on most of the datasets, while the other two performed quite well. However, these larger models still fails to beat the single best DAFT in all cases, except Tweet. This low performance of DAFT models on Tweet is potentially because of the difference in the nature of this dataset. {mdframed}[linecolor=black!0,backgroundcolor=orange!10]These results show that zero-shot performance of DAFT models can vary significantly, making the choice of the DAFT model critical.
3.2.2 Comparison of FT and DA(FT)2
Table 1: Average Relative Improvement of DA(FT)2 compared to FT (Top: 6 sentiment analysis tasks. Bottom: 3 textual similarity tasks): Positive values show DA(FT)2 improvement over FT; larger values imply higher improvements. Please see Table 6 and 7 in Appendix for detailed results. We see that DA(FT)2 significantly improves over FT in few-shot problems.
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
Amazon 75.95 75.25 79.62 65.40 53.51 34.26 0.81
Cornell 67.95 65.61 63.26 60.72 57.77 35.85 8.85
IMDB 74.26 74.64 76.19 66.01 58.20 47.64 0.10
SST2 73.13 72.64 71.81 71.95 71.88 58.58 12.79
Tweet 60.67 54.25 55.53 52.25 41.82 25.92 2.34
Yelp 74.88 81.00 66.81 81.68 46.47 27.86 -0.72
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
MRPC 45.30 43.69 46.97 3.26 32.97 25.40 24.17
QQP 78.40 79.36 94.01 92.48 71.76 52.46 17.84
STSB 49.33 51.92 53.74 57.91 51.06 32.28 3.05
Here we explore the few-shot case when some training data is available from the target domain. The generalized approach with any training data is to fine-tune a model with the data so that the model can perform the domain specific task. The key research question is: should we fine-tune a base model (i.e., FT in Figure 1), or fine-tune a DAFT (i.e., DA(FT)2 in Figure 1) model?  [^4] In our fine-tuning experiments, we found that only updating the weights of the header layer does not guarantee good performance and was easily surpassed by different models designed to perform similar tasks without any fine-tuning (zero-shot). Hence, we chose to do all our performance analysis by focusing on fully fine-tuned models. The performance comparison of updating the weights of the entire model versus only the header layer is provided in the Appendix (Table 6).
Table 1 provides the average performance comparison of the FT and DA(FT)2 models using the Relative Improvement ($RI$) metric when the number of samples to fine-tune FT and DA(FT)2 is varied from $n=2$ to $256$ (The full results for both tasks are in Appendix, Table 6 and 7). The values in the table are the average of $RI$ of DA(FT)2 compared to FT the Roberta base model, when all the DAFT models also have Roberta as the base model. A positive (negative) value of $RI$ means DA(FT)2 is performing better (worse) compared to FT, and the larger (smaller) the value the better (worse) DA(FT)2 is performing. We observe that DA(FT)2 outperforms FT by a large margin when $n$ is within the range of $2$ to $64$. We see only one negative value in the table when $n=256$, which indicates that FT is outperforming DA(FT)2 in most cases. The presence of almost all positive values of $RI$ showcases the importance of using a fine-tuned DAFT instead of FT when the sample size is small.
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]These results indicate that few-shot fine-tuning of DAFT models is generally more beneficial than fine-tuning base models.
The performance of DAFT models showcases the significance of using them for domain adjacent tasks. From a practical standpoint, a major advantage of using DAFT models is that we get them for free. However, as our results show, there is considerable uncertainty about what is the right DAFT model to choose (especially in the zero-shot setting), as the performance can significantly vary between different DAFT models. In the next section, we explore the use of ensemble methods to address this model selection problem.
4 Ensembling of DAFT Models
In this section, we evaluate the performance of the DAFT models using two distinct Ensembling methods and benchmark their performances.
4.1 DAFT-E ${}^{\text{Z}}$ Performance
Figure 4: RI of DAFT ${}^{\text{Z}}$ over DAFT-E ${}^{\text{Z}}$: For each test dataset on the sentiment analysis task, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the average ensemble DAFT-E ${}^{\text{Z}}$ of the 15 DAFT LLMs. Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot prompting on larger LLMs (Bart-large-mnli, Roberta-large-mnli, and Opt-1.3b). The results show that DAFT-E ${}^{\text{Z}}$ usually has significantly better zero-shot performance than DAFT ${}^{\text{Z}}$. However, there are some cases where some specific DAFT models can outperform the average ensemble (leading to $\text{RI}>0$).
The average Ensemble (DAFT-E ${}^{\text{Z}}$) method is a zero-shot method that can be used as an alternative to randomly picking one of the DAFT models. In DAFT-E ${}^{\text{Z}}$, the output probabilities from all the DAFT models are averaged and the decision is taken on that value (max-voting is another feasible approach). To perform average ensembling, we run inference on all the DAFT models. However, since these DAFT models do not need any further training or fine-tuning and inference can be run in parallel, it is feasible to utilize them efficiently with the abundance of computational devices (mostly low end) available today. Figure 4 shows the Relative Improvement of all the 15 DAFT models and the three larger models (as previously discussed in 3.1.2) for sentiment analysis, compared to the average ensemble of the DAFT models, i.e., DAFT-E ${}^{\text{Z}}$.
From Figure 4 we see that very few points have a positive $RI$ value , implying that the average ensemble method (DAFT-E ${}^{\text{Z}}$) is better than most of the DAFT models. Moreover, the performance of the larger LLMs (red squares) are not as good as DAFT-E ${}^{\text{Z}}$. Furthermore, for the Tweet dataset, for which none of the DAFT models performed well and was beaten in performance by two of the larger models (Bart-large-mnli, Roberta-large-mnli), we now have $RI$ values close to zero. Hence, DAFT-E ${}^{\text{Z}}$ has a performance very close to those larger models and shows the importance of using an ensemble even when individual DAFT models cannot perform well. For the text similarity task we see a similar trend in the performance of DAFT-E ${}^{\text{Z}}$ and the $RI$ figure is in the Appendix (Fig: 15). In Section 5, we argue theoretically (Proposition 1) that it is always better to use the average ensemble method over choosing a DAFT model randomly from the set of DAFTs ($\mathcal{M}$).
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]For zero-shot tasks, the ensemble of the DAFT models (DAFT-E ${}^{\text{Z}}$) is on average a better choice than any individual DAFT model.
The average ensemble is therefore a great data-agnostic solution; however, when training data from the target domain is available, we show the data-informed ensembling can be a stronger choice. We thus pursue a weighted ensemble of the DAFT models (DAFT-E) in the following section.
4.2 Comparison of DAFT-E and DA(FT)2
Table 2: Average $RI$ of DA(FT)2 compared to DAFT-E (Top: 6 sentiment analysis tasks. Bottom: 3 text similarity tasks). Negative values indicate that DAFT-E is better than DA(FT)2; larger values imply higher improvements. Table 8 and 9 in the Appendix gives more details. In the very few-shot setting, DAFT-E outperforms DA(FT)2 in almost all cases.
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
Amazon -4.03 -2.48 -2.38 -2.42 -2.66 -2.95 -2.06
Cornell -3.94 -3.60 -3.37 -3.54 -4.13 -3.91 -2.53
IMDB -2.70 -3.41 -3.35 -3.31 -3.04 -2.87 -2.83
SST2 -4.67 -3.32 -3.94 -4.87 -4.11 -4.68 -4.21
Tweet -4.62 -4.91 -5.14 -5.54 -4.00 -3.82 -1.59
Yelp -2.56 -2.76 -2.10 -2.09 -1.68 -2.46 -2.28
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
MRPC -4.91 -3.82 -4.83 -4.96 -2.52 3.12 7.57
QQP 0.22 -0.31 4.70 2.89 5.96 5.14 8.07
STSB 11.30 5.39 22.66 6.18 1.09 2.56 4.78
We define DAFT-E as the weighted ensemble method, where the weights of the ensemble layer are learned using the training data from the target dataset. To perform the weighted ensemble in DAFT-E we utilized two specific regression methods, a) Random forest-based regression ($RF$) and b) SGD-based linear regression ($LR$). The parameters of these methods are discussed in A.5 of the Appendix. We observe the following: (i) $LR$ always has better performance and less variation in performance for smaller sample data, (ii) for higher sample data, i.e., $n$>32, $RF$ usually performs very similar to $LR$, (iii) for sentiment analysis tasks, $LR$ and $RF$ both perform well with similar performance at higher sample data, and (iv for textual similarity tasks, $LR$ tends to show minimal improvement in performance with more samples, whereas $RF$ shows promising results. Based on these observations, and $LR$ being the most lightweight, we show the performance of DAFT-E with $LR$ here and discuss more on the results in the Appendix (Table 8 and 9) and in Section 7 (Limitations).
In the previous section, we showed that DA(FT)2 usually performs better than FT for few-shot learning. Hence, we compare the performance of DAFT-E (weighted ensemble) with DA(FT)2. For a fair comparison, the same amount of data is used when fine-tuning DA(FT)2 and learning the weights of the ensemble layer. Table 2 shows the average RI of DA(FT)2 compared to DAFT-E. Hence, a $RI$ of positive (negative) means the DA(FT)2 performed better (worse, resp.) compared to DAFT-E. From Table 2 it is evident that DAFT-E outperforms DA(FT)2 for sentiment analysis tasks, i.e., all values are negative in the table. On the other hand, for textual similarity DAFT-E usually performs well for smaller data samples (except STSB, see Appendix A.3.2 for discussion), but underperforms on more-data (positive $RI$ values). However, $RF$ performed well with more samples and actually performs better or similarly to DA(FT)2 for $n>16$ (Table 10 in the Appendix). This shows that weighted ensembling is effective, with task-specific variations between $RF$ and $LR$.
Note that updating the ensemble layer depends on the predictions of the DAFT models, and requires solving either (i) a simple linear equation to optimize the weights of the ensemble layer or (ii) running a lightweight $RF$ (max depth = $2$). Thus, only a single inference on each of the DAFT models with the few-shot samples is necessary. In contrast, for fine-tuning a DAFT model, we need to perform back-propagation through the large model to update weights, which is computationally expensive.
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]For few shot tasks, weighted ensembling of DAFT models (DAFT-E) can outperform DA(FT)2, with a suitable ensembling method.
4.3 Overall Comparison
(a) Amazon
(b) Cornell
(c) IMDB
(d) SST2
(e) Tweet
(f) Yelp
Figure 5: Performance comparison of DAFT-E ${}^{\text{Z}}$, DAFT-E, single-best DAFT and FFT. The error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble aggregated over 10 trials. The single best DAFT for different datasets are: Amazon (Roberta - Yelp), Cornell (Roberta - SST2), IMDB (xlnet - amazon), SST2 (xlnet - Cornell), Tweet (xlnet - SST2), Yelp (xlnet - amazon).
For benchmarking purposes, let us assume that we have the entire training data ($D_{T}$) of the target domain, and have a model that is fine-tuned on $D_{T}$ (FFT on $D_{T}$). Let us compare the performance of the ensemble methods with this FFT on different datasets. Figure 5(a)-5(f) show the performance comparison of DAFT-E ${}^{\text{Z}}$ (green dash-dot line), DAFT-E (purple dash-dot line), (single) best DAFT (black dotted line), and FFT on $D_{T}$ (red dotted line) for all the six datasets of the sentiment analysis task. The result for text similarity tasks is provided in the Appendix (Figure 16). From Figure 5, we observe the advantage of using DAFT-E over DAFT-E ${}^{\text{Z}}$ when training data is available. It is interesting to note that DAFT-E has a strong increasing trend in performance (against fine-tuning data size) in most cases, and with the increase of fine-tuning data catches up with either the single best DAFT or FFT for the target domain. The performance of DAFT-E compared to the best solution, i.e., FFT on the target domain, can be bounded theoretically and is given by Proposition 2 in Section 5. Table 3 shows the overall comparison of the computational cost of all the five options that we have discussed in this paper. It should be noted that none of the ensemble methods suffer from the high computational complexity of back-propagation computation; further, using sparse weighting we can reduce the inference cost in DAFT-E compared to DAFT-E ${}^{\text{Z}}$. {mdframed}[linecolor=black!0,backgroundcolor=orange!10]DAFT-E ${}^{\text{Z}}$ and DAFT-E are strong lightweight methods for leveraging already available DAFT models, often matching the single best DAFT model (which is only known post-hoc), and at times matching the best possible performance of full fine-tuning with a large fine-tuning set.
Additional results.
Lastly, DAFT-E ${}^{\text{Z}}$ has shown favorable performance compared to a state-of-the-art zero shot ensemble model, Uniform Soup (Model Soup) Wortsman et al. (2022), and the calculation of LEEP scores Nguyen et al. (2020) on transferability also matches the performance heat map showed in Figure 2. Details of these results with discussion are in Appendix (Tables 11, 12 and Figure 17). Moreover, a table comparing the fundamental concepts of DAFT-E and DAFT ${}^{\text{Z}}$ to the model soup is provided in Table 13
Table 3: Computational cost of methods shown in Figure 1. $n$ denotes the number of few-shot samples. $C_{F}$ and $C_{B}$ denote the computational costs of a forward and backward pass for a LLM. $E$ is the number of fine-tuning epochs. $N$ is the number of DAFT models available to ensemble, $\bar{N}\leq N$ is the number of nonzero weights in the weighted ensemble.
Method Zero-shot Training cost Inference cost
FT $n(C_{F}+C_{B})E$ $C_{F}$
DAFT ${}^{\text{Z}}$ 00 $C_{F}$
DAFT-E ${}^{\text{Z}}$ 00 $N\cdot C_{F}$
DA(FT)2 $n(C_{F}+C_{B})E$ $C_{F}$
DAFT-E $n(C_{F}+E)$ $\bar{N}\cdot C_{F}$
5 Theoretical Analysis
Notations. The dataset of the target domain is $D_{T}$ and the input data to the model is $\mathbf{x}$ and the output data is $\mathbf{y}$, which the model should predict given the input data. There are $N$ number of DAFT models, and a DAFT model can be built using any base model $B_{j}\in\mathcal{B}$, with $j=\{1,2,...,J\}$ and fine-tuning it on dataset $D_{k}\in\mathcal{DA}$ with $k=\{1,2,...,K\}$, where $\mathcal{DA}$ is the set of domain adjacent datasets. The $i^{th}$ DAFT model is given by $M_{i}$ and the corresponding base model and fine-tuning dataset are given by $B_{\kappa(i)}$ and $D_{\nu(i)}$, respectively, i.e., $B_{\kappa(i)}$ was fine-tuned on $D_{\nu(i)}$ dataset to get $M_{i}$. Here, $\kappa(i)$ and $\nu(i)$ are index mapping functions. Also, if we assume that all these DAFT models were created by fine-tuning fully on the given datasets (no fractional fine-tuning), then the total number of DAFT can not be greater than $J\times K$, or $N\leq JK$. Lastly, let $\ell(M(\mathbf{x}),\mathbf{y})$ be the loss in output when model $M$ is used on target dataset input $\mathbf{x}$ and the output is compared with $\mathbf{y}$. Since $\mathbf{x}$ and $\mathbf{y}$ are common for all loss calculation, we can just use $\ell(M)$ to represent $\ell(M(\mathbf{x}),\mathbf{y})$.
5.1 DAFT-E ${}^{\text{Z}}$ versus DAFT ${}^{\text{Z}}$
It is hard to know if a DAFT model is going to perform well or not a priori, and instead of choosing one of the DAFT models in random (DAFT ${}^{\text{Z}}$), one solution could be to use the ensemble of the output results (DAFT-E ${}^{\text{Z}}$) from all the DAFT models.
Proposition 1.
The performance of the average ensemble of DAFT models (DAFT-E ${}^{\text{Z}}$) is no worse than the expected performance obtained from choosing the DAFT models uniformly at random.
Proposition 1 states that in terms of expected performance, using DAFT-E ${}^{\text{Z}}$ is no worse (strictly better, if $\ell$ is assumed to be strictly convex) than picking from the DAFT models uniformly at random.
5.2 DAFT-E versus Optimum
Let us denote $\tilde{M}_{i}$ as the model that uses the base model $B_{\kappa(i)}$ and is fine-tuned on $D_{T}$, i.e., $\tilde{M}_{i}=\Phi(B_{\kappa(i)},D_{T})$. Also, we denote $\tilde{M_{*}}$ is the best performing model when fine-tuned on $D_{T}$ (the base model for $\tilde{M_{*}}$ can be from $\mathcal{B}$ that are used to generate DAFT models or some other large model). The following proposition bounds the difference of loss between the optimum solution and DAFT-E.
Proposition 2.
[…]
$\;+\min_{i\in\mathcal{N}}\left[\mu(\tilde{M}_{i},\tilde{M_{*}})+\rho(D_{\nu(i)},D_{T})\right],$ (1)
where $\mu(\tilde{M}_{i},\tilde{M_{*}})$ is the performance difference between the models $\tilde{M}_{i}$ and $\tilde{M_{*}}$ both fine-tuned on $D_{T}$, and $\rho(D_{\nu(i)},D_{T})$ is an appropriately defined distance measure between $D_{\nu(i)}$ and $D_{T}$.
Corollary 1.
The loss of DAFT-E is larger than the optimum loss by no greater than $min_{i}\rho(D_{\nu(i)},D_{T})$, with the assumption that the base models of the DAFTs can perform as well as the optimum when fine-tuned on the target domain.
Proposition 2 and Corollary 1 (which follows directly from Proposition 2) imply that the performance of DAFT-E depends on the base model on which the DAFT models were trained and the datasets they were fine-tuned on. If the base models (FMs) in the ensemble are nearly as good as the best possible base model (FM) (when the performance of the FFT of the base models are compared), then we can ignore the $\mu()$ term in Equation 1, and the performance of DAFT-E depends only on the dataset that is closest to the target domain. If the DAFT models are generated from a large number of diverse domain adjacent datasets, we expect the distance of the target dataset to the closest DAFT dataset to be small, resulting in DAFT-E performing very close to the optimum.
6 Conclusion
In this paper we introduce the notion of domain adjacent fine-tuned models (DAFT), and investigate how DAFT models can be utilized for inference under limitations on computation time or data required for training a model on a target task. The DAFT models can be used on the target domain (zero-shot) without any fine-tuning. If chosen properly, the performance of DAFT models on the target domain can be close to the optimal performance. However, this performance depends strongly on the choice of the DAFT model, and ensemble methods can be used to address this issue. Two ensembling methods, DAFT-E ${}^{\text{Z}}$ and DAFT-E are explored, and their performances are empirically evaluated and compared against individual DAFT models and other benchmarks involving larger LLMs and base models trained with the full target domain dataset. Our theoretical results support the conclusions from the empirical findings, and provide insights under what conditions DAFT-E can provide near-optimal performance.
7 Limitations
Here in this section we discuss a few limitations and areas of improvement for using DAFT models or ensembles of them.
  • Non domain-adjacent model: DAFT-E or DAFT-E ${}^{\text{Z}}$ will not perform well when none of the available DAFT models are adjacent to the target domain.
  • Sparse weight in Ensemble Layer: Currently DAFT-E ${}^{\text{Z}}$ and DAFT-E needs inference results from all the DAFT models; a sparse weight ensemble may help us select a subset of the DAFT models to have efficient implementation of DAFT-E.
  • Availability of DAFT models: In the ensemble methods discussed, the DAFT models are not trained, rather are used directly from off the shelf in the simplest way. Our argument is that new DAFT models are being generated, and with time DAFT models will be much easier to find. However that is not yet true for some of the tasks being currently investigated in the literature.
  • Run Time: Upon running the DAFT-E and FT in the CPU, we found that the runtime of FT is 6 times more than the inference of the same sample size (Roberta model). Hence, a DAFT-E that ensembles more than 6 DAFT models will be slower than FT, unless the execution of the DAFT models are done in parallel.
Appendix A Appendix
A.1 Datasets and Models
The attributes of the six datasets used for sentiment analysis and the three dataset used for text similarity analysis are given in Table 4 and 5. For ‘SST2’ we had to use validation dataset as test data, because the test data does not have any labels. The direct link to download these datasets are given as follows: https:// huggingface.co /datasets/amazon_polarity, ‘https:// www.cs.cornell.edu /people/pabo/movie-review-data/’, ‘https://huggingface.co/datasets/ imdb’, ‘https:// huggingface.co/ datasets/sst2’, ‘https:// huggingface.co/ datasets/ mteb/tweet_sentiment_extraction’, ‘https:// huggingface.co/ datasets/ yelp_polarity’, ’https://huggingface.co /datasets/nyu-mll/glue’ (when copying the links please remove any spaces from the texts). Figure 6 and 7 shows the performance heat-map for sentiment analysis task when the base models are BERT-based-uncased and xlnet-base-cased respectively. The heat maps in general follows a similar pattern as observed in Figure 2, i.e., Tweet test data cannot be predicted well by any other DAFT models, and Cornell test data is also hard to predict. The observation of DAFT trained on tweet data performing poor is true for bert-based-uncased as well, but not always true for xlnet-base-cased. Another interesting observation is that for xlnet, the DAFT model trained on IMDB data seems to perform quite poorly in a few instances (i.e., on test data of Cornell and Tweet).
Figure 6: Heatmap showing the performance of DAFT models for sentiment analysis task when base model is BERT-based-uncased
Figure 7: Heatmap showing the performance of DAFT models for sentiment analysis task when base model is xlnet-base-cased
A.2 Experimental Settings
The base models were used from huggingface using the ’AutoTokenizer.from_pretrained’, and the webpage with descriptions and examples can be found in ’https://huggingface.co/ transformers/v3.0.2/model_doc/auto.html’. To fine-tune and train these models we used Google Colab platform with the T4 GPU equipped machine. For FT we chose each of the three base models (Roberta, BERT, xlnet) and fine-tuned the models with target data until the loss per epoch did not improve more than 1% or at least a fixed number of data-samples has not been used for training. All these fine-tuned models on the target dataset acted as a suitable candidate for DAFT models. For DA(FT)2 we used any of these DAFT models to fine-tune on another target dataset for few shot training. In the performance comparison of DA(FT)2 we only showed the performance of DA(FT)2 that was fine-tuned on DAFT models having Roberta as its base. For both FT and DA(FT)2 on few shot training, we performed the all the runs five times with five different seeds. For the case of DAFT-E, the weight calculations were done using five different random seeds as well. All the results shown here are the mean values of all those runs.
A.3 Discussion on Performance Anomalies
A.3.1 SST2 on Heatmap
The SST2 dataset obtained from Huggingface did not have any label on its test data, and we had to use the validation data of SST2 as the test data instead. Also, the validation data of SST2 only had $872$ samples. Now, in the heat map shown in Figure 2, the performance of DAFT-SST2 achieved $92\%$ accuracy on the SST2 validation data, whereas, DAFT-Cornell achieved $93.1\%$ accuracy on that same test dataset. This is quite counter intuitive, because we expect DAFT-SST2 to be the best performing one on its own domain (SST2 validation data). We did a thorough simulation analysis to check what might be the reason. Since the DAFT-SST2 was generated using randomly chosen $7,000$ samples from the training dataset of SST2, and we initially extracted the rest of the training data and divided it to different batches of $5,000$ samples. Now, with this batches we checked the performance of DAFT- SST2 and DAFT- Cornell. As expected, we got better performance with DAFT- SST2 compared to DAFT- Cornell with these new batches. With this experiment, we concluded that SST2 validation dataset is, for some unknown reason, more similar to the Cornell training dataset than the SST2 training data, and thus giving a counter intuitive result.
A.3.2 STSB with DAFT-E
The performance of DAFT-E with $LR$ suffers a lot compared to both DA(FT)2 that have been considered here, i.e., DAFT-mrpc and DAFT-qqp. Both of these DAFT models have roberta as their base model, and interestingly these roberta based DAFT models seemed to perform well even at zero shot on STSB dataset. Whereas, the DAFT models based on BERT and xlnet showed considerable less performance. We found this to be the main reason for which DAFT-E performed poorly in comparison to the DA(FT)2 models which already had better performance even without any fine-tuning. To check this, we even checked the DAFT-E performance with DAFT-mrpc and DAFT-qqp with roberta base only, and found that DAFT-E outperformed the DA(FT)2 up to $n=64$.
A.4 Fine-tuning the header or whole model
Figure 10 to 13 show the performance comparison of fine-tuning only the header layer of the base model and fine-tuning the whole model of the base models. It is straightforward to see the superior performance of fine-tuning the whole model even though the base models (i.e., BERT, Roberta) have already gone through extensive training with language data.
Table 4: Dataset sizes and classes (Sentiment Analysis)
Dataset Train data Test data Classes
Amazon 3,600,000 400,000 2
Cornell 7,463 2134 2
IMDB 25,000 25,000 2
SST2 67,349 1821 2
Tweet 12927 3696 3
Yelp 560,000 38,000 2
Table 5: Dataset sizes and classes (Text Similarity)
Dataset Train data Test data Classes
MRPC 3,670 1,730 2
QQP 364,000 40,000 2
STSB 7,500 1,500 2
Figure 8: Performance comparison of FT Header layer to FT whole Model - Amazon dataset.
Figure 9: Performance comparison of FT Header layer to FT whole Model - Cornel dataset.
Figure 10: Performance comparison of FT Header layer to FT whole Model - IMDB dataset.
Figure 11: Performance comparison of FT Header layer to FT whole Model - SST2 dataset.
Figure 12: Performance comparison of FT Header layer to FT whole Model - Tweet dataset.
Figure 13: Performance comparison of FT Header layer to FT whole Model - Yelp dataset.
A.5 Ensemble Layer of DAFT-E
A.5.1 Weight update of $LR$
[…]
$\min_{w_{i}\in\mathbf{w}}\sum_{(X,y)\in D_{T}}\ell\bigg{(}\Big{(}\sum_{i}w_{i}M_{i}(X)\Big{)},y)\bigg{)}+R(\mathbf{w})$
where, $\ell$ is the loss calculated using the output from the DAFT models ($M_{i}(X)$), to the ground truth $y$. Also, the term $R(\mathbf{w})$ can be used to regularize (or control) the number of models ($\bar{N}$) we want to use for our final ensemble. To keep the method simple, we did not use the regularization when implementing DAFT-E.
A.5.2 Parameters of $LR$ and $RF$
For $LR$ we have used SGDRegressor from sklearn.linear_model with maximum iteration of 3. The reason behind this small iteration number is because of the few shot regime. If we had used larger bound on iteration, overfitting might have casued more harm than good. We also used coefficient initialization = $1/N$, where $N$ is the number of DAFT models used, so that at the start DAFT-E gives same weights to all the models. For $RF$ we imported the RandomForestClassifier from sklearn.ensemble, and set the max depth = 2. The smaller max depth was chosen to avoid overfitting.
A.6 Performance of DAFT models and DAFT-E ${}^{\text{Z}}$ for Textual similarity task
Figure 14 and 15 show the Relative Improvement of DAFTs compared to Single Best DAFT and DAFT-E ${}^{\text{Z}}$ respectively for the textual similarity task with three datasets. Similar to the results with sentiment analysis task, we observe the performance deviation among the DAFT models and usual better performance of DAFT-E ${}^{\text{Z}}$.
Figure 14: Relative Improvement of DAFTs compared to the Single Best DAFT for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the single-best DAFT FM (out of the 6). Values less than 0 indicate performance degradation.
Figure 15: Relative Improvement of DAFTs compared to DAFT-E ${}^{\text{Z}}$ for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the DAFT-E ${}^{\text{Z}}$. Values less than 0 indicate performance degradation.
A.7 Proof of Propositions
Proof of Proposition 1
[…]
=\;= $\frac{1}{N}\sum_{i\in\mathcal{N}}\ell(M_{i}(X),y)$
=\;= $\frac{1}{N}\sum_{i\in\mathcal{N}}\ell(\mu_{i},y)\;\;\text{[assuming $M_{i}(X)=\mu_{i}$]}$
[…]
=\;= loss of average ensemble=(DAFT-E $\text{loss of average ensemble}=\;\ell(\text{{DAFT-E}~{}\!${}^{\text{Z}}$}).$ (2)
Note that the derivation assumes $\ell$ to be convex in $\mu_{i}$, the outputs from the DAFT models, which is a common assumption for loss functions. ∎
Proof of Proposition 2
[…]
$=$ loss of weighted ensemble
$=$ $\;\min_{w_{i},i\in\mathcal{N}}\left[\ell\left(\sum_{i}w_{i}M_{i}\right)\right]$
$\leq$ $\;\ell(M_{i}),\;\;\forall i\in\mathcal{N}.$ (3)
Note that the above derivation assumes that the set weights $w_{i},i\in\mathcal{N}$, have been optimized (trained using the fine-tuning dataset from the target domain) to minimize the loss function $\ell\left(\sum_{i}w_{i}M_{i}\right)$.
[…]
$|\ell(M_{1})-\ell(M_{2})|\leq\rho(D_{\nu(1)},D_{\nu(2)}),$ (4)
[…]
$\;\;+\mu(\tilde{M}_{i},\tilde{M}_{*});\;\;\forall i.$ (6)
Since Equation A.7 holds for all $i\in\mathcal{N}$, Equation 1 follows by taking a minimum over all $i$, completing the proof. ∎
Note that Equation 1 bounds the gap between the mimimum loss possible under any model and the loss of DAFT-E. The result is intuitive. The term $\mu()$ tells us that the performance gap depends on how good the base models (corresponding to the DAFT models in the ensemble) are, when compared to the best possible model, when they are all trained on the full target dataset. The term $\rho()$ implies that the performance gap also depends on how closely the datasets used to generate the DAFT models represent the target dataset. Note that the minimum over $i$ is taken on the sum of $\mu()$ and $\rho()$, instead of of each of the two terms individually. However, for an ensemble of DAFT models obtained by fine-tuning a “good” set of FMs with a large number of datasets, the approximation term $\mu()+\rho()$ is expected to be small, as argued in Section 5.2.
Table 6: Relative Improvement of DA(FT)2 compared to FT (Sentiment Analysis)
Dataset    DAFT Name $\;n=2\;$ $\;n=4\;$ $\;n=8\;$ $n=16\;$ $n=32\;$ $n=64\;$ $n=128$ $n=256$
Amazon DAFT- Cornell 80.42 76.28 81.10 67.16 55.45 32.55 0.50 0.48
DAFT- IMDB 78.33 77.79 83.46 68.53 57.12 37.85 2.66 1.92
DAFT- SST2 78.11 75.33 78.67 65.6 54.45 34.45 0.57 0.77
DAFT- Tweet 61.33 67.70 71.34 56.76 47.07 31.19 -1.11 -0.85
DAFT- Yelp 81.59 79.18 83.50 68.95 53.48 35.24 1.42 1.54
Average 75.95 75.25 79.62 65.40 53.51 34.26 0.81 0.77
Cornell DAFT- Amazon 73.60 71.74 68.88 63.6 61.01 38.06 10.25 1.66
DAFT- IMDB 68.43 64.88 64.64 62.43 61.23 38.52 9.58 3.60
DAFT- SST2 74.8 72.11 68.89 66.71 63.54 38.97 9.64 2.67
DAFT- Tweet 53.72 52.43 49.77 48.37 51.15 31.50 6.11 2.23
DAFT- Yelp 69.22 66.87 64.13 62.47 51.87 32.19 8.65 0.3
Average 67.95 65.61 63.26 60.72 57.77 35.85 8.85 2.09
IMDB DAFT- Amazon 81.35 81.84 81.66 70.94 61.86 50.11 2.53 0.65
DAFT- Cornell 77.79 77.09 80.67 66.68 57.4 49.72 2.46 -0.03
DAFT- SST2 76.10 75.28 77.99 64.86 57.45 47.36 -1.33 -0.977
DAFT- Tweet 59.12 60.27 60.49 58.10 53.54 45.20 -0.62 -0.70
DAFT- Yelp 76.93 78.72 80.11 69.47 60.73 45.82 -2.52 0.247
Average 74.26 74.64 76.19 66.01 58.20 47.64 0.10 -0.16
SST2 DAFT- Amazon 76.35 73.84 73.54 70.05 72.62 58.52 11.47 2.03
DAFT- Cornell 82.05 78.97 78.81 80.32 77.56 64.29 16.84 5.41
DAFT- IMDB 71.55 75.20 72.81 74.51 71.62 62.14 14.46 3.85
DAFT- Tweet 62.25 60.66 60.68 62.66 63.46 53.84 10.30 0.92
DAFT- Yelp 73.42 74.51 73.2 72.18 74.17 54.1 10.89 2.75
Average 73.13 72.64 71.81 71.95 71.88 58.58 12.79 2.99
Tweet DAFT- Amazon 60.22 53.93 55.79 41.88 42.67 26.16 1.2 -2.2
DAFT- Cornell 56.37 49.64 50.94 52.62 39.55 23.64 1.03 -1.58
DAFT- IMDB 59.72 53.76 52.58 55.26 45.22 30.19 4.69 1.59
DAFT- SST2 62.23 55.20 57.46 54.1 41.79 26.48 2.29 -1.06
DAFT- Yelp 64.83 58.69 60.85 57.37 39.85 23.15 2.5 -1.6
Average 60.67 54.25 55.53 52.25 41.82 25.92 2.34 -0.97
Yelp DAFT- Amazon 79.81 84.56 71.46 85.73 49.85 26.89 -2.32 -0.41
DAFT- Cornell 74.39 83.52 67.09 81.15 45.62 29.28 -0.65 -0.81
DAFT- IMDB 77.45 83.51 67.84 83.75 47.97 29.62 1.02 0.29
DAFT- SST2 75.4 80.74 66.07 80.69 45.82 26.711 -0.86 -1.48
DAFT- Tweet 67.36 72.68 61.60 77.06 43.10 26.81 -0.79 -1.22
Average 74.88 81.00 66.81 81.68 46.47 27.86 -0.72 -0.73
Table 7: Relative Improvement of DA(FT)2 compared to FT (Text Similarity)
Dataset DAFT Name $\;n=2\;$ $\;n=4\;$ $\;n=8\;$ $n=16\;$ $n=32\;$ $n=64\;$ $n=128$ $n=256$
MRPC DAFT- QQP 34.78 32.60 32.08 -6.49 24.68 18.15 20.21 10.83
DAFT- STSB 55.83 54.79 61.86 13.01 41.26 32.65 28.13 18.20
Average 45.30 43.69 46.97 3.26 32.97 25.40 24.17 14.51
QQP DAFT- MRPC 75.56 78.10 88.81 86.98 66.05 47.51 16.33 4.71
DAFT- STSB 81.24 80.61 99.21 97.98 77.47 57.41 19.36 7.38
Average 78.40 79.36 94.01 92.48 71.76 52.46 17.84 6.05
STSB DAFT- MRPC 46.89 53.02 50.07 51.96 45.97 28.01 1.26 0.64
DAFT- QQP 51.78 50.82 57.42 63.87 56.15 36.55 4.85 1.95
Average 49.33 51.92 53.74 57.91 51.06 32.28 3.05 1.30
Table 8: Relative Improvement of DA(FT)2 compared to DAFT-E (Sentiment Analysis)
Dataset    DAFT Name $\;n=2\;$ $\;n=4\;$ $\;n=8\;$ $n=16\;$ $n=32\;$ $n=64\;$ $n=128$ $n=256$
Amazon DAFT- Cornell -1.60 -1.91 -1.57 -1.38 -1.43 -4.177 -2.35 -1.78
DAFT- IMDB -2.74 -1.07 -0.29 -0.57 -0.38 -0.35 -0.26 -0.37
DAFT- SST2 -2.85 -2.43 -2.89 -2.3 -2.06 -2.80 -2.30 -1.50
DAFT- Tweet -12.01 -6.68 -6.87 -7.51 -6.75 -5.17 -3.92 -3.09
DAFT- Yelp -0.96 -0.30 -0.27 -0.32 -2.683 -2.238 -1.46 -0.753
Average -4.03 -2.48 -2.38 -2.42 -2.66 -2.95 -2.06 -1.50
Cornell DAFT- Amazon -0.71 -0.03 -0.04 -1.81 -2.14 -2.34 -1.27 -3.06
DAFT- IMDB -3.66 -4.02 -2.55 -2.51 -2.03 -2.02 -1.87 -1.21
DAFT- SST2 -0.02 0.187 -0.04 0.05 -0.62 -1.70 -1.82 -2.10
DAFT- Tweet -12.08 -11.27 -11.35 -10.95 -8.15 -6.99 -4.98 -2.52
DAFT- Yelp -3.21 -2.87 -2.85 -2.49 -7.72 -6.50 -2.70 -4.38
Average -3.94 -3.60 -3.37 -3.54 -4.13 -3.91 -2.53 -2.65
IMDB DAFT- Amazon 1.25 0.57 -0.35 -0.43 -0.79 -1.25 -0.48 -0.65
DAFT- Cornell -0.73 -2.05 -0.89 -2.92 -3.53 -1.51 -0.54 -1.32
DAFT- SST2 -1.675 -3.05 -2.36 -3.98 -3.50 -3.06 -4.22 -2.25
DAFT- Tweet -11.16 -11.35 -11.96 -7.92 -5.89 -4.48 -3.53 -1.98
DAFT- Yelp -1.21 -1.25 -1.20 -1.29 -1.49 -4.07 -5.39 -1.05
Average -2.70 -3.41 -3.35 -3.31 -3.04 -2.87 -2.83 -1.45
SST2 DAFT- Amazon -2.89 -2.64 -2.94 -5.91 -3.70 -4.71 -5.33 -4.65
DAFT- Cornell 0.25 0.23 -0.02 -0.24 -0.95 -1.25 -0.77 -1.49
DAFT- IMDB -5.53 -1.88 -3.38 -3.45 -4.26 -2.54 -2.79 -2.95
DAFT- Tweet -10.65 -10.03 -10.16 -10.01 -8.81 -7.53 -6.33 -5.69
DAFT- Yelp -4.50 -2.27 -3.16 -4.74 -2.84 -7.37 -5.82 -3.97
Average -4.67 -3.32 -3.94 -4.87 -4.11 -4.68 -4.21 -3.75
Tweet DAFT- Amazon -4.89 -5.09 -4.98 -11.98 -3.42 -3.64 -2.69 -1.63
DAFT- Cornell -7.17 -7.74 -7.94 -5.31 -5.53 -5.56 -2.86 -1.00
DAFT- IMDB -5.19 -5.21 -6.93 -3.68 -1.69 -0.57 0.67 2.18
DAFT- SST2 -3.70 -4.31 -3.96 - 4.40 -4.02 -3.40 -1.64 -0.48
DAFT- Yelp -2.15 -2.17 -1.89 -2.36 -5.33 -5.94 -1.44 -1.02
Average -4.62 -4.91 -5.14 -5.54 -4.00 -3.82 -1.59 -0.39
Yelp DAFT- Amazon 0.18 -0.84 0.62 0.10 0.60 -3.20 -3.85 -0.98
DAFT- Cornell -2.83 -1.40 -1.93 -2.37 -2.25 -1.37 -2.21 -1.37
DAFT- IMDB -1.13 -1.41 -1.50 -0.97 -0.67 -1.11 -0.57 -0.29
DAFT- SST2 -2.27 -2.90 -2.53 -2.62 -2.11 -3.33 -2.42 -2.04
DAFT- Tweet -6.75 -7.22 -5.16 -4.58 -3.94 -3.26 -2.34 -1.79
Average -2.56 -2.76 -2.10 -2.09 -1.68 -2.46 -2.28 -1.29
Table 9: Relative Improvement of DA(FT)2 compared to DAFT-E- LR (Text Similarity)
Dataset DAFT Name $\;n=2\;$ $\;n=4\;$ $\;n=8\;$ $n=16\;$ $n=32\;$ $n=64\;$ $n=128$ $n=256$
MRPC DAFT- QQP -11.80 -11.25 -14.47 -13.93 -8.60 -2.85 4.15 7.20
DAFT- STSB 1.98 3.60 4.81 4.01 3.56 9.08 11.00 14.33
Average -4.91 -3.82 -4.83 -4.96 -2.52 3.12 7.57 10.77
QQP DAFT- MRPC -1.37 -1.01 1.90 -0.05 2.44 1.73 6.68 7.89
DAFT- STSB 1.82 0.39 7.51 5.83 9.49 8.55 9.46 10.65
Average 0.22 -0.31 4.70 2.89 5.96 5.14 8.07 9.27
STSB DAFT- MRPC 9.48 6.16 19.73 2.18 -2.31 -0.75 2.96 4.29
DAFT- QQP 13.12 4.63 25.59 10.19 4.50 5.87 6.61 5.64
Average 11.30 5.39 22.66 6.18 1.09 2.56 4.78 4.96
Table 10: Relative Improvement of DA(FT)2 compared to DAFT-E- RF (Text Similarity)
Dataset DAFT Name $\;n=2\;$ $\;n=4\;$ $\;n=8\;$ $n=16\;$ $n=32\;$ $n=64\;$ $n=128$ $n=256$
MRPC DAFT- QQP 12.91 -2.02 -10.60 -14.67 -11.38 -7.66 -3.32 -0.07
DAFT- STSB 30.55 14.37 9.56 3.12 0.40 3.67 3.04 6.58
Average 21.73 6.17 -0.52 -5.77 -5.49 -2.00 -0.14 3.26
QQP DAFT- MRPC 5.92 2.20 2.44 -0.88 -1.49 -3.20 1.52 3.59
DAFT- STSB 9.35 3.64 8.08 4.95 5.29 3.29 4.17 6.23
Average 7.64 2.92 5.26 2.03 1.90 0.05 2.85 4.91
STSB DAFT- MRPC 11.08 5.44 -4.39 -5.82 -8.44 -6.09 -2.41 -0.63
DAFT- QQP 14.77 3.92 0.29 1.56 -2.06 0.18 1.05 0.66
Average 12.92 4.68 -2.05 -2.13 -5.25 -2.96 -0.68 0.01
(a) MRPC
(b) QQP
(c) STSB
Figure 16: Performance comparison of DAFT-E ${}^{\text{Z}}$, DAFT-E, single-best DAFT and FFT for Text Similarity task. The error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble, using RF method, aggregated over 10 trials. The single best DAFT for different datasets are: MRPC (Roberta - STSB), QQP (Roberta - STSB), STSB (Roberta - MRPC).
A.8 Comparison with Model Soup
[…]
Dataset Uniform Soup DAFT-E ${}^{\text{Z}}$
Amazon 93.29% 93.72%
Cornell 85.09% 85.37%
IMDB 92.09% 92.28%
SST2 88.07% 89.55%
Tweet 86.09% 87.33%
Yelp 94.67% 95.30%
Table 12: Robustness Comparison of DAFT-E ${}^{\text{Z}}$ and Uniform Soup in terms of the set of models being combined. The results show that as we increase the number of DAFT models, the performance of both Uniform Soup and DAFT-E ${}^{\text{Z}}$ improves, with DAFT-E ${}^{\text{Z}}$ consistently outperforming Uniform Soup in the zero-shot setting. As discussed in Table 11, we are able to add many more models into DAFT-E ${}^{\text{Z}}$ than in Uniform Soup because Uniform Soup requires all models in the ensemble to have matching architectures.
DAFT models combined Uniform Soup DAFT-E ${}^{\text{Z}}$
SST2, Tweet 88.78% 89.09%
SST2, Tweet, Yelp 90.57% 90.86%
SST2, Tweet, Yelp, Cornell 91.19% 91.39%
SST2, Tweet, Yelp, Cornell, Amazon 92.09% 92.28%
Table 13: Comparison of different methods to combine pretrained models in terms of the requirements on the available (DAFT) models that are being combined and the kind of computation necessary. BPF: Method is LLM backpropagation free. SLI: Method needs to perform inference on the target task with a single LLM. NMWA No LLM model weight access is required by the method. MAE: Method can handle multi-architecture ensemble. SFPA: Method requires a single forward pass for few-shot adaptation.
This table highlights that DAFT-E ${}^{\text{Z}}$ and DAFT-E provide lightweight LLM backprop-free model combination schemes (similar to Uniform Soup and Greedy Soup), but DAFT-E ${}^{\text{Z}}$ and DAFT-E do not require access to the model weights, and can combine LLMs with different architectures. Furthermore, the few-shot adaptation in DAFT-E requires only a single forward-pass with the few-shot samples compared to Greedy Soup that requires multiple forward-passes. However, both Uniform and Greedy Soup finally need to perform inference with a single LLM while DAFT-E ${}^{\text{Z}}$ and DAFT-E have to perform inference with multiple LLMs.
Method 0-shot BPF SLI NMWA MAE SFPA
DAFT N/A N/A
Uniform Soup N/A
DAFT-E ${}^{\text{Z}}$ N/A
FT N/A
DA(FT)2 N/A
Greedy Soup
DAFT-E
A.9 Comparison with LEEP
To check the transferability of the DAFT models to a new target domain or task, we checked the LEEP scores between all the 6 datasets of the sentiment analysis (Figure 17). Interestingly, the heat map with the LEEP scores follows closely the heat map that we generated using the performance of the DAFT models (Figure 2). This bolsters the idea of DAFT models and their transferability being proportional to their performance as DAFT models. Moreover, these LEEP scores also tells us that an LEEP scored based ensemble method is possible and needs further investigation.
Figure 17: Heatmap using LEEP scores (Sentiment Analysis)